# Decorrelation Speeds Up Vision Transformers

### v2

@@ Abstract
Keywords: Deep learning, Vision transformers, Efficient AI, Decorrelation
@@ 1 Introduction
Self-attention-based architectures, specifically transformers (vaswani2017attention), have become increasingly popular models both in natural language processing (NLP) and in computer vision. While at first only dominating the NLP domain, transformers have now also become a strong contender in the computer vision domain, competing with or outperforming the previously superior convolutional architectures in tasks such as image classification, segmentation, and detection (carion2020end; dosovitskiy2020image).
@@
The transformer’s self-attention mechanism lets each patch attend to every other patch in an image, resulting in a true global receptive field. In contrast, convolutional neural networks (CNNs) need to stack many local kernels to gradually grow their receptive field over multiple layers (parmar2018image; cordonnier2019relationship; ramachandran2019stand). However, while convolutional layers scale only linearly with image size, self-attention scales quadratically in the number of image patches, incurring steep compute and memory costs at high resolutions (vaswani2017attention; dosovitskiy2020image). This overhead, while tolerable in research experiments, becomes a critical bottleneck in industrial pipelines, where GPU hours are limited and rapid processing of imagery (e.g., for defect segmentation) is essential.
@@
Vision transformers (ViTs) are being widely adopted in real-world inspection and diagnostic pipelines, where annotated data is often scarce but high accuracy is critical. hutten2022vision systematically compared the application of ViT and CNN backbones on railway–freight-car damage assessment imagery and demonstrated that ViTs match or exceed CNN performance in industrial applications with sparsely available data.
@@
Unlike classification, where an image has one label denoting a specific class, segmentation demands pixel-level masks, that classify each pixel as belonging to a particular object. These masks are often difficult to produce, requiring clear boundaries between different objects in the image, and domain-specific experts to perform the annotating. In both industrial-quality control (bovzivc2021mixed) and clinical radiology (bhalgat2018annotation), this is tiresome and costly work and results in only a few hundred annotated examples per needed object class, even though high-resolution imagery is abundant. This stark contrast between abundant unlabelled imagery and scarce segmentation masks indicates a strong need for representation-learning methods that can leverage large amounts of unlabelled imagery and thus reduce the need for extensive pixel-level annotations.
@@
One promising avenue to leverage abundant unlabelled imagery is Masked Autoencoder (MAE) pre-training (he2022masked), which randomly masks a large portion of image patches and trains a ViT to reconstruct the missing pixels. This allows the model to learn rich, generalizable features that transfer well to various downstream tasks when fine-tuned on the limited labelled imagery. However, MAE pre-training on large ViT backbones, such as ViT-Base on ImageNet-1K, can demand tens of GPU-days and hundreds of gigabytes of memory per run, making pre-training on unlabelled data intensive and time consuming (dosovitskiy2020image; he2022masked). On top of the earlier mentioned need for rapid processing of imagery in industrial pipelines, the training of large deep neural networks (DNNs) has also been associated with significant energy consumption and carbon emissions (garcia2019estimation; strubell2020energy; thompson2021deep; de2023growing; luccioni2023estimating; luccioni2024power). Together, these steep GPU requirements, large memory footprints, and non-trivial carbon emissions highlight why MAE pre-training, while beneficial for large amounts of unlabelled imagery, remains impractical for time- and resource-constrained industrial settings.
@@
The high computational cost of training large neural networks has also motivated research into more efficient optimization methods. desjardins2015natural showed that whitening layer activations can approximate natural gradient descent, which improves the conditioning of the optimization problem and accelerates convergence. More recently, ahmad2024correlations argued that correlations in layer inputs distort the gradient direction, and that removing these correlations brings standard gradient descent closer to the natural gradient (amari1998natural). Together, these findings highlight decorrelation as a promising strategy for reducing training cost, making it a natural candidate for addressing the steep computational demands of MAE pre-training.
@@
In this paper we address the impractical training costs of MAE-pretrained ViTs in low-label regimes by applying iterative Decorrelated Backpropagation (DBP) at scale to transformer architectures for the first time. DBP, first introduced by ahmad2022constrained and recently adapted to DNNs by dalm2024efficient, removes correlations in layer inputs to accelerate convergence. We extend and adapt this approach to operate selectively on key blocks of the ViT during MAE pre-training, marking the first large-scale application of DBP in transformer models. This yields: (1) faster pre-training: up to 21.1% wall-time reduction on ImageNet-1K (deng2009imagenet), (2) improved transfer: higher segmentation mIoU when fine-tuning on ADE20K (zhou2017scene), and (3) real-world impact: similar speed and accuracy gains on proprietary industrial datasets.
We validate our approach with MAE pre-training on ImageNet-1K and downstream segmentation on ADE20K, using randomly sampled 10% subsets to mimic data scarcity. Finally, we demonstrate that DBP-accelerated MAE pre-training offers tangible benefits in industrial vision tasks, providing faster iteration and better fine-tuning performance, while also reducing energy use and carbon emissions.
@@ 2.1 Decorrelated Backpropagation
We employ an iterative decorrelation rule first described in ahmad2022constrained. Since decorrelation is applied individually on each layer of a DNN, we first concentrate on describing the decorrelation procedure for one such layer. Given a layer input $\mathbf{x}$ and a decorrelation matrix $\mathbf{R}$, we get the decorrelated input $\mathbf{z}$ as a result of a linear transformation:
$\mathbf{z}=\mathbf{Rx}$ (1)
@@
$\mathbf{R}_{l}\xleftarrow{}\mathbf{R}_{l}-\mathbf{\eta C}_{l}\mathbf{R}_{l}$ (4)
where $\eta$ is a small, constant, and tunable learning rate, and $\mathbf{C}_{l}$ captures the off-diagonal correlations of the layer inputs. Let $X\in\mathbb{R}^{B\times d}$ denote the mini-batch of inputs to layer $l$, with row vectors $x_{l}^{(b)}\in\mathbb{R}^{d}$. We first compute the uncentered covariance matrix
$\mathbf{D}=\frac{1}{B}X^{\top}X\in\mathbb{R}^{d\times d},$ (5)
@@
This removes all self-correlation terms while retaining the pairwise correlations between distinct input dimensions. The objective is to reduce these off-diagonal correlations without encouraging the trivial all-zero solution. For a detailed derivation of this learning rule, we refer to ahmad2022constrained.
@@ 2.2 Application to scaled vision transformers
Our MAE pre-training implementation closely mirrors the ViT-Base recipe of he2022masked. We divide each image into $16\times 16$ patches (as in dosovitskiy2020image), embed them into a 768-dimensional space, and process them through a 12-layer transformer encoder with 12 attention heads per layer. 75% of these patches are randomly masked during training, and a lightweight, 2-layer decoder reconstructs the missing pixels via an $\ell_{2}$ loss. he2022masked report no significant difference in fine-tune performance with decoder depths of 2, 4, or 8 layers. We therefore use a 2-layer decoder for maximal efficiency.
@@
As mentioned in Section 2.1, DBP is applied individually to the layers of a DNN. In our MAE pre-training setup, we extend DBP to the vision transformer architecture by integrating it directly into the encoder. Specifically, we apply DBP to the patch-embedding projection and to all multi-head self-attention and feed-forward MLP layers across the 12 encoder blocks. Figure 1 illustrates this integration. While prior work on vision transformers has leveraged head-decorrelation as a regularizer (e.g. bhattacharyya2023decatt), to our knowledge this is the first work to apply a full decorrelation-based optimization scheme across all encoder layers of a MAE pre-training framework, including patch embeddings and MLPs.
@@
Beyond limiting DBP’s computational overhead to the encoder, we also ensure it remains strictly a pre-training mechanism and does not affect any downstream task. Concretely, for each decorrelated layer we compute the fused weight $\mathbf{\tilde{W}}=\mathbf{WR}$ (see Equation 3) and store $\mathbf{\tilde{W}}$ in place of $(\mathbf{W},\mathbf{R})$. This fusion, performed either at save time or load time, ensures that downstream fine-tuning sees the trained, decorrelated weights directly, without any DBP overhead.
@@ 2.3 Experimental setup
For MAE pre-training we adopt similar parameters used by he2022masked, training for 1000 epochs with a batch size of 4096. We use the AdamW optimizer (loshchilov2017decoupled) with parameters $\beta_{1}=0.9$, $\beta_{2}=0.95$, and a weight decay of 0.05. We optimize for the learning rate, finding an optimal learning rate of $5.0\times 10^{-4}$ for BP-MAE and $1.0\times 10^{-3}$ for DBP-MAE (see Appendix A). We use a cosine decay learning rate scheduler (loshchilov2016sgdr) with a warmup of 40 epochs. The pre-trained models are saved at their point of peak performance, measured in validation loss.
@@
During DBP-MAE pre-training we configure three DBP-specific settings to balance stability, performance, and compute. First, we specify which part of the model to decorrelate, which, as explained in Section 2.2, we set to only the encoder of the model. Next, we modify the number of samples used to measure input correlation during the DBP update. dalm2024efficient found that a subsample of 10% is enough for DBP to achieve the same boost in performance, whilst significantly reducing computational overhead. We empirically found this to be true for our experiments as well, so here too we reduce the number of samples for the DBP update to 10% (see Appendix C). Finally, we optimize for the DBP learning rate (see Equation 4), finding an optimal learning rate of $5.0\times 10^{-4}$ for the DBP update (see Appendix A).
@@
Our architecture setup for fine-tuning uses the pre-trained MAE encoder as the ViT-Base encoder, using UperNet (xiao2018unified) as the decoder. We fine-tune for 125 epochs with a batch size of 16, and use the AdamW optimizer with parameters $\beta_{1}=0.9$, $\beta_{2}=0.999$, and a weight decay of 0.05. We optimize for the learning rate, finding an optimal learning rate of $2.5\times 10^{-5}$, warmed up linearly over the first 500 iterations, then decayed polynomially to zero. Note that we do not apply DBP during fine-tuning. We only differentiate our downstream tasks by their precursor pre-training methods, DBP-MAE or BP-MAE.
@@
We report the pre-training validation loss on the full ImageNet-1K validation set. We validate our fine-tuning performance by reporting the validation loss and the mean Intersection-over-Union (mIoU) on the full ADE20K validation set. Reported wall-clock time was measured as the runtime of the training process during pre-training, excluding validation. Reported carbon emissions were measured as the emissions of the training process during pre-training, excluding validation. Carbon emissions were estimated and tracked using the CodeCarbon python package (benoit_courty_2024_11171501). All models and algorithms were implemented in PyTorch (paszke2019pytorch) and run on a compute cluster using Nvidia A100 or H100 GPUs and AMD EPYC 9334 or Intel Xeon Platinum 8360Y processors. Finally, to complement our benchmark results, we evaluate the method, using the same pre-training setup as described above, on proprietary semiconductor wire bonding imagery, thereby assessing its applicability in real-world industrial use cases. This pre-training dataset is comprised of 7,267 samples, of which 7,220 are used for training and 47 are used for validation. Each sample consists of an image of $2\times 2048\times 2048$ pixels. For the semantic segmentation downstream task, a similar (but labelled) dataset was used, comprised of 5,560 images for training and 788 images for validation. Again, these sets contain images of $2\times 2048\times 2048$ pixels and are annotated for wire, ball, wedge, and epoxy classes. These datasets are provided by ASMPT ALSI B.V. and are not publicly available due to confidentiality agreements.
@@ 3.1 DBP speeds up pre-training convergence
An important thing to remember is that DBP comes with some computational overhead per epoch due to the decorrelation update rule (see Equation 4). Considering this, faster convergence as a function of the number of epochs may not mean that DBP-MAE has more efficient training than BP-MAE. For this reason, we also compare pre-training performance for both regimes as a function of wall-clock time in Figure 3111The wall-clock time measurements in Figure 3 were recorded using 2 NVIDIA H100 GPUs and 2 AMD EPYC 9334 CPUs with a (combined) total core count of 64..
@@ 4 Discussion
In addition to the semiconductor industrial setting, we believe that the DBP-MAE self-supervised technique could also be highly beneficial for medical imaging segmentation. While the medical domain differs substantially from the semiconductor industrial field and presents unique challenges, it similarly suffers from limited datasets and costly, expert-driven annotations. This could be useful particularly for 3D radiological CT and MR scans or large 2D digital pathology slides, where faster pre-training and stronger segmentation performance are critical. Since variants of MAE pre-training have already been applied successfully to medical datasets (zhou2023self; liu2024vis; tang2025hi), accelerating them and further enhancing performance with DBP-MAE represents a promising direction for future research.
@@
Second, the storing and updating of decorrelation matrices required by DBP introduces some additional compute and memory requirements. This is represented by the increased wall-clock time per epoch during training. To reduce this additional computational overhead even further, more effective use could be made of the sparseness structure of the decorrelation matrix. For instance, the large decorrelation matrix could be approximated and updated with the use of low-rank matrix factorization (eckart1936approximation; srebro2003weighted).

### v3

# Kieran Carrigg, Rob van Gastel, Melda Yeghaian, Sander Dalm, Faysal Boughorbel, and Marcel van Gerven

@@ Abstract
keywords:
Deep learning, Vision transformers, Efficient AI, Decorrelation
@@ 1 Introduction
Self-attention-based architectures, specifically transformers [31], have become increasingly popular models both in natural language processing (NLP) and in computer vision. While at first only dominating the NLP domain, transformers have now also become a strong contender in the computer vision domain, competing with or outperforming the previously superior convolutional architectures in tasks such as image classification, segmentation, and detection [7, 14].
@@
The transformer’s self-attention mechanism lets each patch attend to every other patch in an image, resulting in a true global receptive field. In contrast, convolutional neural networks (CNNs) need to stack many local kernels to gradually grow their receptive field over multiple layers [24, 8, 26]. However, while convolutional layers scale only linearly with image size, self-attention scales quadratically in the number of image patches, incurring steep compute and memory costs at high resolutions [31, 14]. This overhead, while tolerable in research experiments, becomes a critical bottleneck in industrial pipelines, where GPU hours are limited and rapid processing of imagery (e.g., for defect segmentation) is essential.
@@
Vision transformers (ViTs) are being widely adopted in real-world inspection and diagnostic pipelines, where annotated data is often scarce but high accuracy is critical. Hütten et al. [18] systematically compared the application of ViT and CNN backbones on railway freight car damage assessment imagery and demonstrated that ViTs match or exceed CNN performance in industrial applications with sparsely available data.
@@
Unlike classification, where an image has one label denoting a specific class, segmentation demands pixel-level masks, that classify each pixel as belonging to a particular object. These masks are often difficult to produce, requiring clear boundaries between different objects in the image, and domain-specific experts to perform the annotating. In both industrial-quality control [6] and clinical radiology [4], this is time intensive and costly work and results in only a few hundred annotated examples per needed object class, even though high-resolution imagery is abundant. This stark contrast between abundant unlabelled imagery and scarce segmentation masks indicates a strong need for representation-learning methods that can leverage large amounts of unlabelled imagery and thus reduce the need for extensive pixel-level annotations.
@@
One promising avenue to leverage abundant unlabelled imagery is Masked Autoencoder (MAE) pre-training [17], which randomly masks a large portion of image patches and trains a ViT to reconstruct the missing pixels. This allows the model to learn rich, generalizable features that transfer well to various downstream tasks when fine-tuned on the limited labelled imagery. However, MAE pre-training on large ViT backbones, such as ViT-Base on ImageNet-1K, can demand tens of GPU-days and hundreds of gigabytes of memory per run, making pre-training on unlabelled data intensive and time-consuming [14, 17]. On top of the earlier mentioned need for rapid processing of imagery in industrial pipelines, the training of large deep neural networks (DNNs) has also been associated with significant energy consumption and carbon emissions [16, 28, 30, 11, 22, 23]. Together, these steep GPU requirements, large memory footprints, and non-trivial carbon emissions highlight why MAE pre-training, while beneficial for large amounts of unlabelled imagery, remains impractical for time- and resource-constrained industrial settings.
@@
The high computational cost of training large neural networks has also motivated research into more efficient optimization methods. Desjardins et al. [13] showed that whitening layer activations can approximate natural gradient descent, which improves the conditioning of the optimization problem and accelerates convergence. More recently, Ahmad [2] argued that correlations in layer inputs distort the gradient direction, and that removing these correlations brings standard gradient descent closer to the natural gradient [3]. Together, these findings highlight decorrelation as a promising strategy for reducing training cost, making it a natural candidate for addressing the steep computational demands of MAE pre-training.
@@
In this paper we address the impractical training costs of MAE-pretrained ViTs in low-label regimes by applying iterative Decorrelated Backpropagation (DBP) at scale to transformer architectures for the first time. DBP, first introduced by Ahmad et al. [1] and recently adapted to DNNs by Dalm et al. [10], removes correlations in layer inputs to accelerate convergence. We extend and adapt this approach to operate selectively on key blocks of the ViT during MAE pre-training, marking the first large-scale application of DBP in transformer models. This yields: (1) faster pre-training: up to 21.1% wall-time reduction on ImageNet-1K [12], (2) improved transfer: higher segmentation mIoU when fine-tuning on ADE20K [33], and (3) real-world impact: similar speed and accuracy gains on proprietary industrial datasets.
The remainder of this paper is organized as follows. Section 2 describes the DBP algorithm, its implementation for Vision Transformers, and the MAE pre-training and ViT fine-tuning experimental setups. Section 3 presents results for MAE pre-training on ImageNet-1K and downstream segmentation on ADE20K, using randomly sampled 10% subsets to simulate data scarcity, as well as DBP-accelerated MAE pre-training on proprietary industrial datasets. Section 4 discusses the implications and limitations of these results, outlines directions for future research, and concludes.
@@ 2.1 Decorrelated Backpropagation
We employ an iterative decorrelation rule first described in Ahmad et al. [1]. Since decorrelation is applied individually on each layer of a DNN, we first concentrate on describing the decorrelation procedure for one such layer. Given a layer input $\mathbf{x}$ and a decorrelation matrix $\mathbf{R}$, we get the decorrelated input $\mathbf{z}$ as a result of a linear transformation:
$\mathbf{z}=\mathbf{Rx}.$ (1)
@@
$\mathbf{R}_{l}\xleftarrow{}\mathbf{R}_{l}-\eta\mathbf{C}_{l}\mathbf{R}_{l}$ (4)
where $\eta$ is a small, constant, and tunable learning rate, and $\mathbf{C}_{l}$ captures the off-diagonal correlations of the layer inputs. Let $\mathbf{X}\in\mathbb{R}^{B\times d}$ denote the mini-batch of inputs to layer $l$, with row vectors $\mathbf{x}_{l}^{(b)}\in\mathbb{R}^{d}$. We first compute the uncentered covariance matrix
$\mathbf{D}=\frac{1}{B}\mathbf{X}^{\top}\mathbf{X}\in\mathbb{R}^{d\times d}$ (5)
@@
This removes all self-correlation terms while retaining the pairwise correlations between distinct input dimensions. The objective is to reduce these off-diagonal correlations without encouraging the trivial all-zero solution. For a detailed derivation of this learning rule, we refer to Ahmad et al. [1].
@@ 2.2 Application to scaled vision transformers
Our MAE pre-training implementation closely mirrors the ViT-Base recipe of He et al. [17]. We divide each image into $16\times 16$ patches (as in Dosovitskiy et al. [14]), embed them into a 768-dimensional space, and process them through a 12-layer transformer encoder with 12 attention heads per layer. 75% of these patches are randomly masked during training, and a lightweight, 2-layer decoder reconstructs the missing pixels via an $\ell_{2}$ loss. He et al. [17] report no significant difference in fine-tune performance with decoder depths of 2, 4, or 8 layers. We therefore use a 2-layer decoder for maximal efficiency.
@@
As mentioned in Section 2.1, DBP is applied individually to the layers of a DNN. In our MAE pre-training setup, we extend DBP to the vision transformer architecture by integrating it directly into the encoder. Specifically, we apply DBP to the patch-embedding projection and to all multi-head self-attention and feed-forward MLP layers across the 12 encoder blocks. Figure 1 illustrates this integration. While prior work on vision transformers has leveraged head-decorrelation as a regularizer (e.g. Bhattacharyya et al. [5]), to our knowledge this is the first work to apply a full decorrelation-based optimization scheme across all encoder layers of a MAE pre-training framework, including patch embeddings and MLPs.
@@
Beyond limiting DBP’s computational overhead to the encoder, we also ensure it remains strictly a pre-training mechanism and does not affect any downstream task. Concretely, for each decorrelated layer we compute the fused weight $\mathbf{\tilde{W}}=\mathbf{WR}$ as implied by Eq. (3), and store $\mathbf{\tilde{W}}$ in place of $(\mathbf{W},\mathbf{R})$. This fusion, performed either at save time or load time, ensures that downstream fine-tuning sees the trained, decorrelated weights directly, without any DBP overhead.
@@ 2.3 Experimental setup
For MAE pre-training we adopt similar parameters used by He et al. [17], training for 1000 epochs with a batch size of 4096. We use the AdamW optimizer [21] with parameters $\beta_{1}=0.9$, $\beta_{2}=0.95$, and a weight decay of 0.05. We optimize for the learning rate, finding an optimal learning rate of $5.0\times 10^{-4}$ for BP-MAE and $1.0\times 10^{-3}$ for DBP-MAE (see Appendix A). We use a cosine decay learning rate scheduler [20] with a warmup of 40 epochs. The pre-trained models are saved at their point of peak performance, measured in validation loss.
@@
During DBP-MAE pre-training we configure three DBP-specific settings to balance stability, performance, and compute. First, we specify which part of the model to decorrelate, which, as explained in Section 2.2, we set to only the encoder of the model. Next, we modify the number of samples used to measure input correlation during the DBP update. Dalm et al. [10] found that a subsample of 10% is enough for DBP to achieve the same boost in performance, whilst significantly reducing computational overhead. We empirically found this to be true for our experiments as well, so here too we reduce the number of samples for the DBP update to 10% (see Appendix C). Finally, we optimize for the DBP learning rate in Eq. (4), finding an optimal learning rate of $5.0\times 10^{-4}$ for the DBP update (see Appendix A).
@@
Our architecture setup for fine-tuning uses the pre-trained MAE encoder as the ViT-Base encoder, using UperNet [32] as the decoder. We fine-tune for 125 epochs with a batch size of 16, and use the AdamW optimizer with parameters $\beta_{1}=0.9$, $\beta_{2}=0.999$, and a weight decay of 0.05. We optimize for the learning rate, finding an optimal learning rate of $2.5\times 10^{-5}$, warmed up linearly over the first 500 iterations, then decayed polynomially to zero. Note that we do not apply DBP during fine-tuning. We only differentiate our downstream tasks by their precursor pre-training methods, DBP-MAE or BP-MAE.
@@
We report the pre-training validation loss on the full ImageNet-1K validation set. We validate our fine-tuning performance by reporting the validation loss and the mean Intersection-over-Union (mIoU) on the full ADE20K validation set. Reported wall-clock time was measured as the runtime of the training process during pre-training, excluding validation. Reported carbon emissions were measured as the emissions of the training process during pre-training, excluding validation. Carbon emissions were estimated and tracked using the CodeCarbon python package [9]. All models and algorithms were implemented in PyTorch [25] and run on a compute cluster using Nvidia A100 or H100 GPUs and AMD EPYC 9334 or Intel Xeon Platinum 8360Y processors. Finally, to complement our benchmark results, we evaluate the method, using the same pre-training setup as described above, on proprietary semiconductor wire bonding imagery, thereby assessing its applicability in real-world industrial use cases. This pre-training dataset is comprised of 7,267 samples, of which 7,220 are used for training and 47 are used for validation. Each sample consists of an image of $2\times 2048\times 2048$ pixels. For the semantic segmentation downstream task, a similar (but labelled) dataset was used, comprised of 5,560 images for training and 788 images for validation. Again, these sets contain images of $2\times 2048\times 2048$ pixels and are annotated for wire, ball, wedge, and epoxy classes. These datasets are provided by ASMPT ALSI B.V. and are not publicly available due to confidentiality agreements.
@@ 3.1 DBP speeds up pre-training convergence
An important thing to remember is that DBP comes with some computational overhead per epoch due to the decorrelation update rule in Eq. (4). Considering this, faster convergence as a function of the number of epochs may not mean that DBP-MAE has more efficient training than BP-MAE. For this reason, we also compare pre-training performance for both regimes as a function of wall-clock time in Figure 3111The wall-clock time measurements in Figure 3 were recorded using 2 NVIDIA H100 GPUs and 2 AMD EPYC 9334 CPUs with a (combined) total core count of 64..
@@ 4 Discussion
In addition to the semiconductor industrial setting, we believe that the DBP-MAE self-supervised technique could also be highly beneficial for medical imaging segmentation. While the medical domain differs substantially from the semiconductor industrial field and presents unique challenges, it similarly suffers from limited datasets and costly, expert-driven annotations. This could be useful particularly for 3D radiological CT and MR scans or large 2D digital pathology slides, where faster pre-training and stronger segmentation performance are critical. Since variants of MAE pre-training have already been applied successfully to medical datasets [34, 19, 29], accelerating them and further enhancing performance with DBP-MAE represents a promising direction for future research.
@@
Second, the storing and updating of decorrelation matrices required by DBP introduces some additional compute and memory requirements. This is represented by the increased wall-clock time per epoch during training. To reduce this additional computational overhead even further, more effective use could be made of the sparseness structure of the decorrelation matrix. For instance, the large decorrelation matrix could be approximated and updated with the use of low-rank matrix factorization [15, 27].